OpenStudy (anonymous):

the amount of paper recycled is directly proportional to the number of trees that recycling can save. if recycling 4,000 pounds of paper saves 34 trees how many trees are saved are saved when 6,000 pounds of paper is recycled??

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} \textit{recycled paper}&\textit{saved trees}\\ \hline\\ 4000&34\\ 6000&x \end{array}\implies \bf \cfrac{4000}{6000}=\cfrac{34}{x}\)

OpenStudy (anonymous):

what

OpenStudy (jdoe0001):

solve for "x"

OpenStudy (anonymous):

If paper recycled 4000 pounds then tees saved=34 .........................6000 .................................=(34/4000)*6000

OpenStudy (anonymous):

ohh thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so the answer is 24000000

OpenStudy (anonymous):

34 trees / 4000 lbs = x trees / 6000 lbs Cross-multiplying and solving for x = 51 trees

OpenStudy (anonymous):

\[\frac{ 34 }{ 4000 }*6000=\frac{ 34*6 }{ 4 }=51\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw